what is the probability that a market inspector will discover greater than five violation of the public health code

1
Expert's answer
2022-02-10T04:19:26-0500

Let the random variable "X" represent the number of violations of the public health code.

The probability that a market inspector will discover greater than 5 violations of health code is, "p(X\gt5)=\displaystyle \sum^{\infin}_{x=6}p(x)" . This is equivalent to "p(X\gt5)=1-p(x\le5)=1-\{p(0)+p(1)+p(2)+p(3)+p(4)+p(5)\}"